OpenStudy (anonymous):

evaluate -13+|3.1-4.2|

OpenStudy (anonymous):

its not equal to anything?

OpenStudy (anonymous):

my options are -14.1, -11.9 or -5.7

OpenStudy (anonymous):

\[-13+\left| -1.1\right|\] -13+(-).(-1.1)=-13+1.1=11.9

OpenStudy (anonymous):

-13+(-).(-1.1)=-13+1.1=-11.9 sorry

OpenStudy (anonymous):

i agree with cinar

OpenStudy (anonymous):

he/she is correct